# Small solutions to inhomogeneous and homogeneous quadratic congruences modulo prime powers

### v3

@@ 1. Introduction and statement of results
since the squares of the variables are much smaller than the modulus $q$. In this case, the number of solutions in question is zero if $Q$ is definite and satisfies an asymptotic of the form $\sim C_{p}N^{n-2}\log N$ if $Q$ is definite, where $C_{p}$ is a suitable constant depending on $p$.
@@
This relates to the following result by Heath-Brown for the case of odd square-free moduli, which was established by a different method.
Theorem 1.4 (Theorem 1 in [11]).
Fix $\varepsilon>0$. Let $q$ be odd and square-free and $Q$ be any ternary quadratic form with integer coefficients whose determinant is coprime to $q$. Then the congruence $Q(x_{1},x_{2},x_{3})\equiv 0\bmod{q}$ has a solution $(x_{1},x_{2},x_{3})\in\mathbb{Z}\setminus\{0,0,0\}$ satisfying $\max\{|x_{1}|,|x_{2}|,|x_{3}|\}\ll q^{5/8+\varepsilon}$.
In [8] and [9], Heath-Brown established, among other results, that for integral quadratic forms $Q$ in more than 3 variables, the smallest non-zero solution of the congruence $Q(x_{1},...,x_{n})\equiv 0\bmod{q}$ is bounded by $O\left(q^{1/2+\varepsilon}\right)$ if $q$ has at most two prime factors. For general moduli $q$, he established a bound of $O\left(q^{1/2+3/n^{2}+\varepsilon}\right)$.
Inhomogeneous diagonal forms of degree $k\geq 2$ were considered by Cochrane, Ostergaard and Spencer in [4] and [5]. In particular, they proved that for any positive integers $k,n$ with $n>3(k^{2}+k+2)/2$, prime $p$, and integers $\lambda_{1},...,\lambda_{n+1}$ with $(\lambda_{1}\cdots\lambda_{n},p)=1$, there exists a solution $(x_{1},...,x_{n})$ to the congruence
$\lambda_{1}x_{1}^{k}+\cdots+\lambda_{n}x_{n}^{k}\equiv\lambda_{n+1}\bmod{p}$
with $1\leq x_{i}\ll_{k}p^{1/k}$ for i=1,,ni=1,...,n. The above condition on $n$ becomes $n>12$ if $k=2$.
The goals of the present article are two-fold. Firstly, we aim to prove a variant of Theorem 1.2 for inhomogeneous quadratic forms modulo odd prime powers $q=p^{m}$. Secondly, we aim to show that Theorem 1.2 remains valid for homogeneous diagonal forms in more than 3 variables if the coefficients are allowed to vary with the modulus. That is, if $n>3$, the said result holds for arbitrary coefficients whenever $N\geq q^{1/2+\varepsilon}$. In contrast, for ternary forms, it appears difficult to improve the exponent $11/18$ in Theorem 1.3 to $1/2$. We emphasize that we establish asymptotic formulae, whereas the above-mentioned results in [4], [5], [8], [9] and [11] concern the existence of small solutions. Moreover, our conditions on the variables are more restrictive than in the above-mentioned results by Heath-Brown and make the problem interesting for prime powers (see the discussion in [1, section 1]). Our first main result in this article is the following.
@@
Let $\Phi:\mathbb{R}\to\mathbb{R}_{\geq 0}$ be a Schwartz class function. Then as $m\to\infty$, we have the asymptotic formula
@@
with $q=p^{m}$, provided that $N\geq q^{1/2+\varepsilon}$.
@@
(i) In Theorem 1.5, we have assumed the quadratic form $\lambda_{1}x_{1}^{2}+\cdots+\lambda_{n}x_{n}^{2}$ to be positive definite. It should be possible to drop this condition. However, this would depend on an asymptotic formula for the weigthed number of representations by indefinite quadratic forms which is currently not at our disposal. We will say more about this in Remark 2.6 below.
(ii) The set $\mathcal{A}(Q)$ in Theorem 1.6 above is non-empty as a consequence of [14][parts (iii) and (iv) of Theorem 6 on page 37]. We note that under the conditions of our Theorem 1.6, $\varepsilon=1=(-1,-1)$, in Serre’s notation.
(iii) The set $\mathcal{B}(Q)$ in Theorem 1.5 turns out to be non-empty as well, but we will establish this by a different argument below. We note that in Theorem 1.5, the condition $(x_{1},...,x_{n})\not\equiv(0,....0)\bmod{p}$ in Theorem 1.6 is replaced by the stronger condition $(x_{1}\cdots x_{n},p)=1$. This stronger condition helps us in our later calculations because it results in taking differences of Gauss sums which often cancel each other, as seen in subsection 3.3.1 below. To see that $\mathcal{B}(Q)$ is non-empty, we note that if $p\geq 5$, then each of the sets
@@
$S_{2}:=\left\{\lambda-\lambda_{n-1}x_{n-1}^{2}-\lambda_{n}x_{n}^{2}\in\mathbb{F}_{p}:(x_{n-1},x_{n})\in\left(\mathbb{F}_{p}^{\times}\right)^{2}\right\}$
@@
(iv) One can prove a result of the same strength as in Theorem 1.6 with the condition $(x_{1},...,x_{n})\not=(0,...,0)\bmod{p}$ replaced by the condition $(x_{1}\cdots x_{n},p)=1$ in Theorem 1.5. To this end, one may follow the same steps as in the proof of Theorem 1.5 up to equation (3.6) and then evaluate the sum $K_{n}(-A,-\lambda_{n+1},p^{m-r})$ for $\lambda_{n+1}=0$. This leads to a similar bound for the error term $U$ as in (4.2). However, we decided to be in line with Theorem 1.2 and directly use our expressions for this error term in [2].
@@
where $\Lambda$ is coprime to $p$, $\sqrt{k\Lambda}$ denotes one of the two square roots of $k\Lambda$ modulo $p^{s}$, if existent (otherwise, the summand is omitted), and $\tau_{n}(k)$ denotes a weigthed number of integral solutions $(x_{1},...,x_{n})$ to a Diophantine equation of the form
@@
We point out that a possible alternative approach to the above problems consists of writing the congruence $Q({\bf x})\equiv 0\bmod{q}$ as an equation $Q({\bf x})=lq$, detecting weighted solutions of this equation directly using the circle method and summing up over $l$ in the relevant range. However, isolating main terms in the simple forms as they appear in our Theorems 1.5 and 1.6 above then requires much more work. In contrast, the approach in this paper yields these main terms in a natural and easy way, and for the error terms we obtain essentially optimal bounds.
Data availability statement: This manuscript has no associated data.
Conflict of interest statement: The authors have no conflicts of interest to declare. All co-authors have seen and agree with the contents of the manuscript and there is no financial interest to report. We certify that the submission is original work and is not under review at any other publication.
Acknowledgements: The authors thank the Ramakrishna Mission Vivekananda Educational and Research Institute for excellent working conditions. The second- and third-named authors both would like to thank CSIR, Govt. of India for financial support in the form of a Senior Research Fellowship under file numbers 09/934(0015)/2019-EMR-I and 09/934(0016)/2019-EMR-I, respectively.
@@ 2.3. Results on Gauss sums
Let $a,b\in\mathbb{Z}$ and $c\in\mathbb{N}$.
(i) If $(a,c)$ divides $b$, then
@@
(v) If $c$ is any natural number and $(a,c)=1$, then
$|G(a,0,c)|\leq\sqrt{c}.$
@@
These properties of quadratic Gauss sums can be found in [3][Chapter 1]. ∎
@@ 2.4. Results on Kloosterman and Salié sums
This is a part of [6][Theorem 3.1(iii)]. ∎
@@
See [12][Page 322, Exercise 1] and [12][Page 323, Lemma 12.4] for parts (i) and (ii). To prove part (iii) we use the Proposition 2.3 as follows. The Kloosterman sum $K_{0}(a,b,c)$ can be expressed as
@@
2.5. Weigthed representations by quadratic forms
We shall use the following result on weigthed numbers of respresentations by quadratic forms, which is a variant of [13][Theorem 1.1].
@@ 2.5. Weigthed representations by quadratic forms
Fix an odd prime $p$ and a homogeneous quadratic form $F^{(0)}(x_{1},...,x_{n})\in\mathbb{Z}[x_{1},...,x_{n}]$ in $n\geq 4$ variables which is non-singular modulo $p$ and positive definite over $\mathbb{R}$. Fix a Schwartz class function $\Omega:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$. Suppose that $k\in\mathbb{N}$ and $P\geq 1$. For ${\bf x}=(x_{1},...,x_{n})\in\mathbb{R}^{n}$ set
@@
Our proposition above can be established along the same lines as [13][Theorem 1.1] using the Kloosterman refinement of the circle method, following the treatment in [12][section 20.4]. Below we explain in which points it differs from [13][Theorem 1.1] and why this doesn’t create extra difficulties.
(i) Our weight function $w$ is a product of Schwartz class functions, whereas in [13][Theorem 1.1], it is a bump function. (ii) In [13][Theorem 1.1], the precise dependence of the error term on the quadratic form was worked out. Here we assume that this form is fixed, and hence the error term simplifies. (iii) There is an additional summation condition $(x_{1}\cdots x_{n},p)=1$ on the left-hand side of (2.2).
Point (i) is not a problem because [13][Theorem 1.1] can be, more generally, established for Schwartz class weight functions. This requires a modification of the principle on nonstationary phase (see [13][section 5.1]) for Schwartz class functions which is possible due to the rapid decay of $w({\bf x})$ and its partial derivatives of all orders. Point (ii) just makes things easier. Point (iii) is reflected in the singular series, where we also have an extra summation condition of the same form in the definition of the summands $a_{q}(k)$. We note that here the modulus for the sum over $\bf x$} is extended from $q$ to $pq$, which is compensated by an extra factor of $p^{-n}$. ∎
Remark 2.6.
In [10][Theorem 4], a somewhat similar result on weigthed numbers of respresentations by quadratic forms was established using Heath-Brown’s variant of the circle method. This result has the advantages that it also covers indefinite forms and comes with an error term of smaller size $k^{n/4-1/2+\varepsilon}$, which is roughly the square root of the main term. However, it is restricted to the particular case when $P=k^{1/2}$. If this result could be extended to the situation of general $P$ and Schwartz class weight functions, similarly as in Proposition 2.5 above, then the condition of positive-definiteness in our Theorem 1.5 could be dropped.
@@
(i) There is a positive constant $M$ only depending on $w$ such that $|\sigma_{\infty,P}(k)|\leq M$ if $k\in\mathbb{N}$ and $P\geq 1$.
@@
where $A>0$ is arbitrary, the implied constant only depending on $F^{(0)}$, $w$ and $A$. This is because $w({\bf x})$ is a product of Schwartz class functions and $F^{(0)}({\bf x})$ is positive definite. It follows that
@@
and hence $\sigma_{\infty,P}(k)=O(1)$, establishing the claim in part (i).
@@
$f_{1}(t)=\left(\frac{n}{2}-1\right)\cdot t^{-1}\sigma_{\infty,P}(t)$
and hence
$f_{1}(t)\ll t^{-1}$ (2.4)
using part (i). Calculating $\frac{d}{dt}w\left(t^{1/2}{\bf x}\right)$ using the definition of $w$, we obtain

### v4

@@ 1. Introduction and statement of results
The existence of small solutions of congruences of the form
$F(x_{1},...,x_{n})\equiv\lambda\bmod{q},$ (1.1)
where $F$ is a form in $n$ variables, has received a lot of attention. Of particular interest is the case of diagonal forms
$F(x_{1},...,x_{n})=\lambda_{1}x_{1}^{k}+\cdots+\lambda_{n}x_{n}^{k}.$
A very general result in this direction was obtained by Cochrane, Ostergaard and Spencer [3] who established the existence of a solution to (1.1) for this case in any cube of sidelength $q^{1/k+\varepsilon}$ if $q$ is a prime, $k\geq 2$, $0<\varepsilon<1/(k(k-1))$ and $n>(k-1)/\varepsilon$. Their paper [3] also provides a history of results on this problem. The homogeneous case when $\lambda\equiv 0\bmod{q}$ is generally easier to handle and has therefore received more attention than the inhomogeneous case. Particularly strong results have been established for homogeneous quadratic congruences ($\lambda=0$ and $k=2$) in at least three variables ($n\geq 3$) by Schinzel, Schlickewei and Schmidt [14], Heath-Brown [8, 9, 11] and Hakimi [6]. In the present article, we are interested in asymptotic formulae for the number of small solutions of quadratic congruences (rather than just existence results), with a special emphasis on the more difficult inhomogeneous case. We will confine ourselves to prime power moduli $q=p^{m}$. This allows us to evaluate Kloosterman and Salié sums coming up in our method explicitly and use Hensel-type arguments. Below we review recent research on asymptotic formulae for the number of small solutions of homogeneous quadratic congruences to prime power moduli.
@@
since the squares of the variables are much smaller than the modulus $q$. In this case, the number of solutions in question is zero if $Q$ is definite and satisfies an asymptotic of the form $\sim C_{p}N^{n-2}\log N$ if $Q$ is indefinite, where $C_{p}$ is a suitable constant depending on $p$.
@@
This should be compared to Heath-Brown’s result in [11, Theorem 1] for composite moduli $q$ and general quadratic forms in three variables, which establishes the existence of a non-trivial solution of height $\max\{|x_{1}|,|x_{2}|,|x_{3}|\}\ll q^{5/8+\varepsilon}$.
The goals of the present article are two-fold. Firstly, we aim to prove a variant of Theorem 1.2 for inhomogeneous quadratic forms. This constitutes the larger part of this article. Secondly, we aim to show that Theorem 1.2 remains valid for homogeneous diagonal forms in more than 3 variables if the coefficients are allowed to vary with the modulus. That is, if $n>3$, the said result holds for arbitrary coefficients whenever $N\geq q^{1/2+\varepsilon}$. In contrast, for ternary forms, it appears difficult to improve the exponent $11/18$ in Theorem 1.3 to $1/2$.
Our first main result in this article is the following.
@@
Let $\Phi:\mathbb{R}\to\mathbb{R}_{\geq 0}$ be a Schwartz class function whose Fourier transform is a bump function. Then as $m\to\infty$, we have the asymptotic formula
@@
We point out that if $k=2$ and $q$ is a prime, then the result of Cochrane, Ostergaard and Spencer mentioned at the beginning of this paper requires the number of variables $n$ to be larger than $1/\varepsilon$ to establish the existence of a non-trivial solution of height $\max\{|x_{1}|,...,|x_{n}|\}\leq q^{1/2+\varepsilon}$ to the congruence in question.
@@
(i) In Theorem 1.4, we have assumed the quadratic form $\lambda_{1}x_{1}^{2}+\cdots+\lambda_{n}x_{n}^{2}$ to be positive definite. It should be possible to drop this condition. However, this would depend on an asymptotic formula for the weighted number of representations by indefinite quadratic forms which is currently not at our disposal.
(ii) Moreover, in Theorem 1.4, we have assumed the Fourier transform of our Schwartz class weight function $\Phi$ to be a bump function, which is a commonly used assumption in analytic number theory. We note that we also assumed $\Phi$ to be real-valued and non-negative on $\mathbb{R}$. It is easy to construct a function $\Phi$ with these properties: Take any bump function $f:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$ and form the convolution $g=f\ast f^{\ast}$, where $f^{\ast}(x):=f(-x)$. This function $g$ is a bump function as well and its Fourier transform equals $|\hat{f}|^{2}$. The function $\Phi=|\hat{f}|^{2}$ is in the Schwartz class and non-negative, and its Fourier transform satisfies $\hat{\Phi}(x)=g(-x)=g(x)$ and hence is a bump function. To remove the condition that $\hat{\Phi}$ is a bump function from Theorem 1.4, we would need to be able to establish a more general version of [13, Theorem 1.1] (resp., Proposition 2.5 below) for Schwartz class instead of bump functions. This would require to prove an analogue of [13, Theorem 5.13] for Schwartz class weight functions, which is a non-trivial task. We leave this to future research.
(iii) The set $\mathcal{A}(Q)$ in Theorem 1.5 above is non-empty as a consequence of [15, parts (iii) and (iv) of Theorem 6 on page 37]. We note that under the conditions of our Theorem 1.5, $\varepsilon=1=(-1,-1)$, in Serre’s notation.
(iv) The set $\mathcal{B}(Q)$ in Theorem 1.4 turns out to be non-empty as well, but we will establish this by a different argument below. We note that in Theorem 1.4, the condition $(x_{1},...,x_{n})\not\equiv(0,....0)\bmod{p}$ in Theorem 1.5 is replaced by the stronger condition $(x_{1}\cdots x_{n},p)=1$. This stronger condition helps us in our later calculations because it results in taking differences of Gauss sums which often cancel each other, as seen in subsection 3.3.1 below. To see that $\mathcal{B}(Q)$ is non-empty, we note that if $p\geq 5$, then each of the sets
@@
$S_{2}:=\left\{\lambda_{n+1}-\lambda_{n-1}x_{n-1}^{2}-\lambda_{n}x_{n}^{2}\in\mathbb{F}_{p}:(x_{n-1},x_{n})\in\left(\mathbb{F}_{p}^{\times}\right)^{2}\right\}$
@@
where $\Lambda$ is coprime to $p$, $\sqrt{k\Lambda}$ denotes one of the two square roots of $k\Lambda$ modulo $p^{s}$, if existent (otherwise, the summand is omitted), and $\tau_{n}(k)$ denotes a weighted number of integral solutions $(x_{1},...,x_{n})$ to a Diophantine equation of the form
@@
We point out that a possible alternative approach to the above problems consists of writing the congruence $Q({\bf x})\equiv 0\bmod{q}$ as an equation $Q({\bf x})=lq$, detecting weighted solutions of this equation directly using the circle method and summing up over $l$ in the relevant range. However, isolating main terms in the simple forms as they appear in our Theorems 1.4 and 1.5 above then requires much more work. In contrast, the approach in this paper yields these main terms in a natural and easy way, and for the error terms we obtain essentially optimal bounds.
Acknowledgements. The authors would like to thank the anonymous referees for their valuable comments. They would also like to thank the Ramakrishna Mission Vivekananda Educational and Research Institute for providing excellent working conditions. The second-named author would like to thank CSIR, Govt. of India for financial support in the form of a Junior Research Fellowship under file number 09/934(0015)/2019-EMR-I.
Data availability statement. No data are associated to this article.
Conflict of interest statement. The authors declare no conflict of interest regarding this article.
@@ 2.3. Results on Gauss sums
Let $a,b\in\mathbb{Z}$ and $c\in\mathbb{N}$. Then we have the following.
(i) If $(a,c)|b$, then
@@
(v) If $(a,c)=1$, then
$|G(a,0,c)|\leq 2\sqrt{c}.$
@@
(i) Suppose that $(a,c)=d$ and $d|b$. Set $\tilde{a}:=a/d$, $\tilde{b}=b/d$ and $\tilde{c}=c/d$. Then we have
[…]
establishing part (i).
(ii) Suppose that $(a,c)=d$ and $d\nmid b$. Set $\tilde{a}:=a/d$ and $\tilde{c}:=c/d$. Then we have
$G(a,b,c)=\sum_{n=0}^{c-1}e_{c}(an^{2}+bn)=\sum\limits_{n=0}^{c-1}e_{\tilde{c}}\left(\tilde{a}n^{2}\right)e_{c}\left(bn\right)=\sum\limits_{n=0}^{\tilde{c}-1}e_{\tilde{c}}\left(\tilde{a}n^{2}\right)\sum\limits_{\begin{subarray}{c}m=0\\ m\equiv n\bmod{\tilde{c}}\end{subarray}}^{c-1}e_{c}(bm)=0$
since
$\sum\limits_{\begin{subarray}{c}m=0\\ m\equiv n\bmod{\tilde{c}}\end{subarray}}^{c-1}e_{c}(bm)=\sum\limits_{k=0}^{d-1}e_{c}(b(k\tilde{c}+n))=e_{c}(bn)\sum\limits_{k=0}^{d-1}e_{d}(bk)=0,$
establishing part (ii).
(iii) Suppose that $(a,c)=1$ and $c$ is odd. Then quadratic completion gives
$G(a,b,c)=\sum_{n=0}^{c-1}e_{c}(an^{2}+bn)=e\left(-\frac{\overline{4}ab^{2}}{c}\right)\sum_{n=0}^{c-1}e_{c}\left(a(n+\overline{2a}b)^{2}\right)=e\left(-\frac{\overline{4}ab^{2}}{c}\right)\sum\limits_{n=0}^{c-1}e_{c}(an^{2})=e\left(-\frac{\overline{4}ab^{2}}{c}\right)G(a,0,c),$
establishing part (iii).
Parts (iv) and (v) are consequences of [5, Lemmas 7.12-15.]. ∎
@@ 2.4. Results on Kloosterman and Salié sums
This is a part of [4, Theorem 3.1(iii)]. ∎
@@
See [12, Page 322, Exercise 1] and [12, Page 323, Lemma 12.4] for parts (i) and (ii). To prove part (iii) we use the Proposition 2.3 as follows. The Kloosterman sum $K_{0}(a,b,c)$ can be expressed as
@@
2.5. Weighted representations by quadratic forms
We shall use the following result on weighted numbers of representations by quadratic forms, which is a variant of [13, Theorem 1.1].
@@ 2.5. Weighted representations by quadratic forms
Fix an odd prime $p$ and a homogeneous quadratic form $F^{(0)}(x_{1},...,x_{n})\in\mathbb{Z}[x_{1},...,x_{n}]$ in $n\geq 4$ variables which is non-singular modulo $p$ and positive definite over $\mathbb{R}$. Fix a bump function $\Omega:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$. Suppose that $k\in\mathbb{N}$ and $P\geq 1$. For ${\bf x}=(x_{1},...,x_{n})\in\mathbb{R}^{n}$ set
@@
Our proposition above is a slightly modified version of [13, Theorem 1.1]. These two results differ in the following points.
(i) In [13, Theorem 1.1], the precise dependence of the error term on the quadratic form was worked out. Here we assume that this form is fixed, and hence the error term simplifies.
(ii) There is an additional summation condition $(x_{1}\cdots x_{n},p)=1$ on the left-hand side of (2.2). This reflects in the singular series, where we also have an extra summation condition of the same form in the definition of the summands $a_{q}(k)$. We note that here the modulus for the sum over $\bf x$} is extended from $q$ to $pq$, which is compensated by an extra factor of $p^{-n}$.
Proposition 2.5 can be established along the same lines as [13, Theorem 1.1] using the Kloosterman refinement of the circle method, following the treatment in [12, section 20.4]. ∎
@@
(i) There is a positive constant $M$ only depending on $w$ such that $|\sigma_{\infty,P}(t)|\leq M$ if $t\geq 1$ and $P\geq 1$.
@@
where $A>0$ is arbitrary, the implied constant only depending on $F^{(0)}$, $w$ and $A$. This is because $w({\bf x})$ is a product of bump functions and $F^{(0)}({\bf x})$ is positive definite. It follows that
@@
and hence $\sigma_{\infty,P}(t)=O(1)$, establishing the claim in part (i).
@@
$\begin{split}f_{1}(t)\ll&t^{n/2-2}\lim\limits_{\varepsilon\rightarrow 0^{+}}\frac{1}{2\epsilon}\int\limits_{|F^{(0)}({\bf x})-1/P^{2}|<\epsilon}w(t^{1/2}{\bf x}){\rm d}{\bf x}\\ =&t^{n/2-1}\lim\limits_{\varepsilon\rightarrow 0^{+}}\frac{1}{2\epsilon}\int\limits_{|F^{(0)}({\bf x})-1/P^{2}|<\epsilon/t}w(t^{1/2}{\bf x}){\rm d}{\bf x}\\ =&t^{-1}\lim\limits_{\varepsilon\rightarrow 0^{+}}\frac{1}{2\epsilon}\int\limits_{|F^{(0)}(t^{-1/2}{\bf x})-1/P^{2}|<\epsilon/t}w({\bf x}){\rm d}{\bf x}\\ =&t^{-1}\lim\limits_{\varepsilon\rightarrow 0^{+}}\frac{1}{2\epsilon}\int\limits_{|F^{(0)}({\bf x})-t/P^{2}|<\epsilon}w({\bf x}){\rm d}{\bf x}\\ =&t^{-1}\sigma_{\infty,P}(t)\\ \ll&t^{-1},\end{split}$ (2.4)
using part (i) for the last line. Calculating $\frac{d}{dt}w\left(t^{1/2}{\bf x}\right)$ using the definition of $w$, we obtain